A proton is constant to rotate in a circle of radius 125 pm around the fixed point given the rotational energy to the classical energy at 25 degree c calculate value of ml.
1
Expert's answer
2021-09-15T02:41:22-0400

Mass of a proton = 1.6726 x 10^-27 kg, classical average energy=1/2kBT, where kBT is Boltzman constant = 1.30 x 10^ -23 J K^-1, and T is the temperature.) 

Ml = 125/25=5ml